HUBBLE SPACE TELESCOPE DAILY REPORT #2039 PERIOD COVERED: 0000Z (UTC) 01/13/98 - 0000Z (UTC) 01/14/98 Daily Status Report as of 014/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed NIC/1/2/3 7456 (Water in Protoplanetary Disks) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the protoplanetary disk around the star HD19305 to search for water in its disks. This survey will be the first to take a detailed look at the gas phase properties of pre-main-sequence disks at AU distances. The observations were completed as scheduled, and no anomalies were reported. 1.2 Completed Seven Sets of WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed STIS/CCD 7802 (Daily Darks to Update Acquisition Bad Pixel Table) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations to obtain daily darks to update the acquisition bad pixel table. These darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed Seven Sets of NIC/1/3 7813 (NIC/3 HDF Darks) The Near Infrared Camera (Cameras One and Three) was used to make internal dark calibration images which will be used when processing the Hubble Deep Field images during the first NICMOS Camera Three campaign. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed NIC/1/2/3 7322 (The IMF Below 1 Msun in Young Clusters: 1.9micron Water Band Imaging of IC348) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of clusters GL-569AB and LHS-3003. This study will be the first to obtain spectral types for a young cluster population complete to K=16, and thereby quantify, to unprecedentedly high precision, the IMF below 1M{sun} in young clusters. The observations were executed as scheduled, and no anomalies were reported. 1.6 Completed Five Sets of WF/PC-2/NIC/1/2/3 7257 (An Unbiased Redshift Survey of Ultra-Faint Galaxies) The WF/PC-2 and the Near Infrared Camera (Cameras One, Two and Three) were used to make observations of the ultra-faint galaxy A851-(Positions 1, 2, 3 and 4) as part of an unbiassed redshift survey. Near-infrared observations of these types of objects will provide a robust estimate of the luminosity density in the distant universe, an important constraint on theories of galaxy formation. The observations were completed as planned, and no problems were reported. 1.7 Completed Two Sets of NIC/1/2/3 7239 (Surface Compositions of Kuiper Belt Objects) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the Kuiper Belt object Pholus in order to study its surface composition. The surfaces of Kuiper Belt Objects contain a record of formative and dynamical processes in the outer solar system. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 11 Successful : 11 Following a successful acquisition that started at 014/013916Z, the V3 observer bias exhibited erratic values around "0", with a mean value of 1.97a-s/hr and a standard deviation of 8.53 a-s/hr. Both values did not meet the COP criteria. As a result, a bad V3 bias was saved at 014/0219Z (11.52 a-s/hr) that had no effect on the following acquisition. HSTAR 6313 was written. Scheduled Reacquisitions : 5 Successful : 5 2.2 FHST Updates: Scheduled : 19 Successful : 17 As documented in HSTAR 6314, the updates at 014/011836Z and 014/025524Z failed with no impact to science. Both were at approximately the same attitude. 2.3 Operations Notes: The -E SPA was re-connected at 013/1331Z. HSTARs 6311 and 6312 were filed by the STScI to document: 1) a target not being in the FOV for NICMOS target acquisition test proposal 7808 and 2) a failure to locate the coronographic hole during the same test. Both events occurred on day 363 of last year. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH